# Static for Dynamic: Towards a Deeper Understanding of Dynamic Facial Expressions Using Static Expression Data

### v2

@@ Abstract
Index Terms:
Dynamic facial expression recognition, mixture of experts, self-supervised learning, vision transformer.
@@ I Introduction
Based on these observations, we propose a novel framework, Static-for-Dynamic (S4D), which enhances the understanding of dynamic facial expressions by fully leveraging static expression data at both the input- and label- levels. To the best of our knowledge, we represent the first systematic exploration and analysis of the inherent correlations between SFER and DFER tasks, integrating both modalities in terms of input data (i.e., facial images and videos) and emotion labels. The S4D framework employs a unified dual-modal learning approach to seamlessly merge SFER and DFER tasks, thereby advancing DFER performance. Specifically, S4D comprises two key stages: Dual-Modal Pre-Training and Joint Fine-Tuning. During the pre-training stage, we utilize a shared Vision Transformer (ViT) [17] encoder-decoder architecture combined with Masked Autoencoders (MAE) [18] to perform masked modeling on both image and video modalities. This approach allows the model to efficiently learn generalizable spatiotemporal representations from both static and dynamic facial data. In the fine-tuning stage, the pre-trained encoder is jointly optimized using both FER image and video datasets with a multi-task learning setup, facilitating cross-modal interaction of emotional information. However, as highlighted in previous works [19, 20], directly employing multi-task learning in related FER tasks may result in a negative transfer. We also observed similar performance degradation when applying straightforward multi-task learning to SFER and DFER tasks. To address this issue, we introduce a novel Mixture of Adapter Experts (MoAE) module, which operates in parallel with the original Feed-Forward Network (FFN) in the ViT layers. This innovative design enables the model to simultaneously capture task-agnostic knowledge through the FFN and task-specific features via the MoAE, thereby preventing negative transfer and enhancing emotional information interaction. By integrating the MoAE module within the unified dual-modal learning framework, the S4D framework effectively integrates complementary information from static facial expression data, leading to a deeper understanding of dynamic facial expressions and significant improvements in DFER performance. The code and model are publicly available here [^1] https://github.com/MSA-LMC/S4D.
@@ II-A Dynamic Facial Expression Recognition
The evolution of Dynamic Facial Expression Recognition (DFER) has undergone a significant paradigm shift, transitioning from traditional laboratory-based approaches relying on handcrafted features to advanced deep learning-based methods capable of handling in-the-wild scenarios. Recent DFER approaches can be broadly categorized into four distinct groups. The first group, end-to-end supervised learning, initially employed 3D CNNs [22] such as C3D [23], R(2+1)D [24], and I3D-RGB [25] to learn spatiotemporal features from raw videos. This approach evolved to a two-stage process, combining 2D CNNs [22] for spatial feature extraction with RNNs RNNs [26] or LSTMs [27, 10, 28] for temporal modeling. More recently, Transformer-based architectures [29], such as those explored in works like[30, 31, 32], have become dominant in this category. The second group leverages advanced vision-language models, particularly CLIP [33], to advance DFER techniques. Prominent methods such as CLIPER [34], DFER-CLIP [5], and A3lign-DFER [35] capitalize on CLIP’s semantic understanding, enhancing expression recognition by bridging visual and textual modalities. The third group employs self-supervised learning to exploit large amounts of unlabeled facial video data [6, 36, 37]. Notably, MAE-DFER [6] utilizes a local-global interaction Transformer encoder for masked reconstruction, improving task-specific learning. The fourth and final group explores transfer learning from static knowledge to enhance DFER performance [7, 16]. For instance, AEN [16] combines multi-level semantic features from SFER models with emotion-guided loss functions, while S2D [7] adapts a pre-trained SFER model to DFER via Temporal Modeling Adapters. Unlike these methods, which focus on single-modal or task-specific learning, our S4D framework integrates both static and dynamic facial expression data through a unified dual-modal learning approach, offering a more comprehensive and promising solution for DFER.
@@ IV-A Datasets
Method Sample Strategies Backbone DFEW FERV39k MAFW
@@ IV-B Implementation Details
We adopt ViT-B/16 [17] as the backbone for S4D and VideoMAE [8, 6] as the baseline. To reduce background noise, we crop a central 160$\times$160 patch from each 224$\times$224 frame in the VoxCeleb2 dataset, following [6]. For pre-training, we sample 16 frames per clip with a temporal stride of 4, use a patch size of $2\times 16\times 16$. Static images are resized to 160$\times$160 and temporally replicated to match the patch size. We apply random masking at a ratio of 95% for videos and 90% for images, consistent with [43]. Training is performed using the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$, an over batch size $N_{\mathrm{bs}}$ of 384, and a base learning rate $lr_{\mathrm{base}}$ of 1.6e-3, with weight decay set to 0.05. The learning rate is scaled linearly based on batch size: $lr=lr_{\mathrm{base}}\times\frac{N_{\mathrm{bs}}}{512}$. Pre-training is conducted on the VoxCeleb2 and AffectNet-8 datasets for 100 epochs, with a cosine learning rate scheduler.
For fine-tuning, we replace the latter half of the ViT layers with MoAE layers and initialize the model using the pre-trained ViT encoder weights. All input frames and images are resized to $160\times 160$. The batch size is 32 for the DFER datasets and 64 for the SFER datasets, with a learning rate of $4\times 10^{-5}$. We sample 16 frames per clip with a temporal stride of 4 across all datasets, except for FERV39K, which uses a stride of 1 due to its unique properties. In the MoAE module, we set $k=2$ and the number of experts $n=8$. During joint fine-tuning, the proportion of the SFER dataset in each epoch is empirically set to 50%. We optimize the model using AdamW for 100 epochs. During inference, following [6, 7], we uniformly sample two clips per video and averaging their predictions for DFER tasks. Unless otherwise specified, all results are obtained using the S4D (ViT+MoAE) model with the best-performing weights.
@@ IV-C Comparison with the State of the Art
Additionally, Table III presents a detailed analysis of S4D’s category-specific performance on the DFEW dataset, demonstrating its strong overall accuracy. Specifically, S4D achieves the highest accuracy in happy (94.76%), surprise (67.26%), disgust (27.59%), and fear (44.57%), with a notable 9.66% improvement in disgust over the baseline method VideoMAE [6]. It also ranks second in sad (79.04%), neutral (74.60%), and angry (79.80%). Notably, S4D excels in traditionally challenging categories (surprise, fear, and disgust), where data imbalance is prevalent, due to its effective dual-modal pre-training and joint fine-tuning strategy. This approach enables S4D to construct representations with reduced bias while effectively leveraging the complementary information from SFER data through unified dual-modal learning, resulting in more robust emotion recognition across all categories. Although the improvement in disgust accuracy is noteworthy, the relatively low absolute accuracy (27.59%) underscores the inherent difficulty in recognizing this subtle emotion and suggests potential data imbalance within the DFEW dataset. Further investigation into techniques for mitigating data imbalance may yield additional improvements in this category.
@@ IV-D1 Impact of Dual-Modal Pre-training
Table IV analyzes the impact of various pre-training settings on DFER performance. All pre-training methods result in significant improvements in both UAR and WAR compared to the no pre-training setting, emphasizing the importance of pre-training in DFER. Specifically, pre-training on single-modal video data outperforms image-based pre-training, as it captures crucial temporal information. However, image-based pre-training also proves beneficial by enabling the model to learn fine-grained facial features that complement the temporal representations derived from video data. Dual-modal pre-training, which combines image (AffectNet) and video (VoxCeleb2) datasets (Experiment #4), achieves the best performance, with a UAR/WAR of 43.41%/53.65% on FERV39K and 70.58%/79.37% UAR/WAR on DFEW. This confirms that dual-modal pre-training effectively captures both temporal information and fine-grained facial details, leading to robust representation capabilities. Experiment #5, which utilized homogeneous image and video data from VoxCeleb2 for dual-modal pre-training, exhibited slightly lower performance than Experiment #4. This suggests that leveraging diverse datasets and modalities for pre-training can yield richer and more beneficial feature representations. Overall, dual-modal pre-training significantly enhances DFER performance, surpassing both single-modal and no pre-training approaches.
TABLE IV: Comparison of the proposed dual-modal pre-training with other pre-training settings. Dual-modal pre-training on AffectNet and VoxCeleb2 significantly improves DFER performance and outperforms single-modal pre-training settings with less pre-training cost. AFC: AffectNet; VC2: VoxCeleb2.
# Pre-Training Setting (images + videos) Pre-Training Cost (TFLOPs) FERV39K DFEW
@@
1 Without pre-training 0 27.67 39.56 40.96 52.65
2 Image only (AFC) $8.8\times 10^{4}$ 36.10 47.59 53.80 66.82
3 Video only (VC2) $1.8\times 10^{6}$ 42.92 52.58 67.71 76.33
4 Dual-modal (AFC + VC2) $\bm{1.4\times 10^{6}}$ 43.41 53.65 70.58 79.37
5 Dual-modal (VC2 + VC2) $1.4\times 10^{6}$ 42.36 52.94 68.35 78.72
@@ IV-D2 Pre-Training Cost Analysis
As shown in Table IV, the proposed dual-modal pre-training (Experiment #4 and #5) demonstrates superior efficiency and effectiveness, significantly reducing computational costs compared to video-only pre-training (Experiment #3). Specifically, the integration of image data reduces the pre-training cost by approximately 22% (from 1.8 to 1.4 $\times$ 106 TFLOPs), achieved through aggressive masking ratios of 90% for images and 95% for videos, minimizing redundant computations. This approach not only optimizes efficiency but also enhances performance, yielding notable improvements on the FERV39K (+0.49% UAR, +1.07% WAR) and DFEW (+2.87% UAR, +3.04% WAR) benchmarks. In contrast, image-only pre-training (Experiment #2) incurs a minimal cost (8.8 $\times$ 104 TFLOPs) but shows limited effectiveness for DFER tasks. Similarly, the no pre-training setting (Experiment #1) incurs no pre-training cost but severely degrades performance. These findings underscore the critical importance of capturing temporal dynamics for DFER, which are absent in image-only pre-training or no pre-training strategies.
@@
# Model Setting Params (M) FERV39K DFEW
@@ IV-D3 Evaluation of the MoAE Module
To evaluate MoAE’s effectiveness in mitigating negative transfer between SFER and DFER tasks, we conducted comprehensive comparisons with other multi-task learning (MTL) methods. As shown in Table V, the vanilla ViT model achieves a performance of 42.68%/53.08% UAR/WAR on FERV39K and 68.69%/77.01% UAR/WAR on DFEW dataset. When SFER data is incorporated through simple multi-task learning (simply adding multiple task-specific heads to ViT, Experiment #7), we observe minimal improvement or even slight performance degradation compared to the vanilla ViT. This phenomenon suggests that straightforward multi-task adaptation may fail to capture task relationships and lead to negative transfer. In contrast, our proposed S4D (with MoAE) demonstrates significant improvements across all metrics while maintaining reasonable parameter efficiency. Specifically, S4D achieves the best performance with 43.41%/53.65% UAR/WAR on FERV39K and 70.58%/79.37% UAR/WAR on DFEW, validating its effectiveness in learning both task-specific knowledge and generic features. To further investigate the architectural benefits, we compared S4D with ViT-MoE [48], a classic MoE architecture that replaces the standard FFN with FFN experts. Despite utilizing substantially more parameters (115M/285M vs. 90M/101M for testing/training), ViT-MoE exhibits poorer performance. These results emphasize that S4D’s superior performance stems from its architectural innovations rather than mere parameter scaling, highlighting its parameter efficiency and effectiveness in learning complementary knowledge between SFER and DFER tasks while mitigating negative transfer.
TABLE VI: Ablation studies on the joint fine-tuning strategy. The joint fine-tuning strategy significantly enhances performance, demonstrating the complementary benefits of integrating static and dynamic expression data. *: Produced by the epoch at 43 for optimal SFER performance.
Fine-Tuning Setting (data mixture) FERV39K AffectNet-7
@@ IV-D5 Contribution of Proposed Components
We conducted ablation studies to evaluate the impact of each component in our S4D framework: dual-modal pre-training (D.P.), MoAE, and joint fine-tuning (J.F.). Results from Table VII indicate that D.P. improved UAR/WAR by 0.10%/0.74% on FERV39K and 4.03%/0.85% on DFEW over the baseline (our fine-tuned version of VideoMAE on the pre-trained weights from [6]), underscoring its role in enhancing spatiotemporal representations. However, MoAE alone led to a slight performance decrease on all metrics, suggesting it may require a joint fine-tuning strategy to unlock its task-specific capabilities fully. The optimal configuration, integrating all components, achieved a UAR/WAR of 43.41%/53.65% on FERV39K and 70.58%/79.37% on DFEW datasets, highlighting the synergistic benefits of our S4D framework in refining DFER performance through discriminative feature learning.

### v3

@@ I Introduction
Based on these observations, we propose a novel framework, Static-for-Dynamic (S4D), which enhances the understanding of dynamic facial expressions by fully leveraging static expression data at both the input- and label- levels. To the best of our knowledge, we represent the first systematic exploration and analysis of the inherent correlations between SFER and DFER tasks, integrating both modalities in terms of input data (i.e., facial images and videos) and emotion labels. The S4D framework employs a unified dual-modal learning approach to seamlessly merge SFER and DFER tasks, thereby advancing DFER performance. Specifically, S4D comprises two key stages: Dual-Modal Pre-Training and Joint Fine-Tuning. During the pre-training stage, we utilize a shared Vision Transformer (ViT) [17] encoder-decoder architecture combined with Masked Autoencoders (MAE) [18] to perform masked modeling on both image and video modalities. This approach allows the model to efficiently learn generalizable spatiotemporal representations from both static and dynamic facial data. In the fine-tuning stage, the pre-trained encoder is jointly optimized using both FER image and video datasets with a multi-task learning setup, facilitating cross-modal interaction of emotional information. However, as highlighted in previous works [19, 20], directly employing multi-task learning in related FER tasks may result in a negative transfer. We also observed similar performance degradation when applying straightforward multi-task learning to SFER and DFER tasks. To address this issue, we introduce a novel Mixture of Adapter Experts (MoAE) module, which operates in parallel with the original Feed-Forward Network (FFN) in the ViT layers. This innovative design enables the model to simultaneously capture task-agnostic knowledge through the FFN and task-specific features via the MoAE, thereby preventing negative transfer and enhancing emotional information interaction. By integrating the MoAE module within the unified dual-modal learning framework, the S4D framework effectively integrates complementary information from static facial expression data, leading to a deeper understanding of dynamic facial expressions and significant improvements in DFER performance. The code and models are publicly available at https://github.com/MSA-LMC/S4D.
@@ II-A Dynamic Facial Expression Recognition
The evolution of Dynamic Facial Expression Recognition (DFER) has shifted from traditional handcrafted feature-based methods to advanced deep learning approaches for in-the-wild scenarios. Recent DFER approaches can be broadly categorized into four distinct groups. First, end-to-end supervised learning initially used 3D CNNs like C3D [22], R(2+1)D [23], and I3D-RGB [24] for spatiotemporal feature learning, later combining 2D CNNs [25] with RNNs or LSTMs [26, 27] for temporal modeling, and more recently adopting Transformer-based architectures [28, 29, 30]. Second, CLIPER [31], DFER-CLIP [5] and A3lign-DFER [32] utilize vision-language models, notably CLIP [33] to improve DFER performance by incorporating semantic understanding. Thid,  [6, 34, 35] employ self-supervised learning to exploit large amounts of unlabeled facial video data. Notably, MAE-DFER [6] utilizes a local-global interaction Transformer encoder for masked reconstruction, improving task-specific learning. Finally, recent works explore transfer learning from static knowledge to enhance DFER performance [7, 16]. For instance, AEN [16] combines multi-level semantic features from SFER models with emotion-guided loss functions, while S2D [7] adapts a pre-trained SFER model to DFER via Temporal Modeling Adapters. Unlike these methods, which focus on single-modal or task-specific learning, our S4D framework integrates both static and dynamic facial expression data through a unified dual-modal learning approach, offering a more comprehensive and promising solution for DFER.
@@ IV-A Datasets
Method
Sample
Strategies
Backbone DFEW FERV39k MAFW
@@ IV-B Implementation Details
We adopt ViT-B/16 [17] as the backbone for both our S4D, following VideoMAE [8] and MAE-DFER [6]. MAE-DFER initialized VideoMAE model with the weights pretrained exclusively on the VoxCeleb2 dataset and fine-tuned on the DFER dataset, serving as a fair and comparable baseline. To reduce background noise, we crop a central $\times160$ patch from each $\times224$ frame in the VoxCeleb2 dataset, following [6]. For pre-training, we sample 16 frames per clip with a temporal stride of 4, use a patch size of $2\times 16\times 16$. Static images are resized to $\times160$ and temporally replicated to match the patch size. We apply random masking at a ratio of 95% for videos and 90% for images, consistent with [41]. Training is performed using the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$, an over batch size $N_{\mathrm{bs}}$ of 384, and a base learning rate $lr_{\mathrm{base}}$ of 1.6e-3, with weight decay set to 0.05. The learning rate is scaled linearly based on batch size: $lr=lr_{\mathrm{base}}\times\frac{N_{\mathrm{bs}}}{512}$. Pre-training is conducted on the VoxCeleb2 and AffectNet-8 datasets for 100 epochs, with a cosine learning rate scheduler.
For fine-tuning, we replace the latter half of the ViT layers with MoAE layers and initialize the model using the pre-trained ViT encoder weights. We employ a full fine-tuning strategy, where all parameters of the model are updated during training. All input frames and images are resized to $160\times 160$. The batch size is 32 for the DFER datasets and 64 for the SFER datasets, with a learning rate of $4\times 10^{-5}$. We sample 16 frames per clip with a temporal stride of 4 across all datasets, except for FERV39K, which uses a stride of 1 due to its unique properties. In the MoAE module, we set $k=2$ and the number of experts $n=8$. During joint fine-tuning, the proportion of the SFER dataset in each epoch is empirically set to 50%. We optimize the model using AdamW for 100 epochs. During inference, following [6, 7], we uniformly sample two clips per video and averaging their predictions for DFER tasks. Unless otherwise specified, all results are obtained using the S4D (ViT+MoAE) model with the best-performing weights.
@@ IV-C Comparison with the State of the Art
Additionally, Table III presents a detailed analysis of S4D’s category-specific performance on the DFEW dataset, clearly illustrating its substantial gains across both majority and minority emotion classes. Specifically, S4D achieves the highest accuracy in happy (94.76%), surprise (67.26%), disgust (27.59%), and fear (44.57%), with a notable 9.66% improvement in disgust over the baseline method VideoMAE [6]. It also ranks second in sad (79.04%), neutral (74.60%), and angry (79.80%). Importantly, S4D demonstrates particularly strong improvements in traditionally underrepresented and challenging categories such as surprise, fear, and disgust, which are often affected by data imbalance. These improvements are largely attributed to S4D’s dual-modal pre-training and joint fine-tuning framework, which effectively exploits the complementary information from SFER data through unified dual-modal learning, resulting in more robust emotion recognition across all categories. Although the improvement in disgust accuracy is noteworthy, the relatively low absolute accuracy (27.59%) underscores the inherent difficulty in recognizing this subtle emotion and suggests potential data imbalance within the DFEW dataset. Further investigation into techniques for mitigating data imbalance may yield additional improvements in this category.
@@ IV-D1 Impact of Dual-Modal Pre-training
Table IV analyzes the impact of various pre-training settings on DFER performance. All pre-training methods result in significant improvements in both UAR and WAR compared to the no pre-training setting, emphasizing the importance of pre-training in DFER. Specifically, pre-training on single-modal video data (lines 3 and 4) outperforms image-based pre-training, as it captures crucial temporal information. However, image-based pre-training also proves beneficial by enabling the model to learn fine-grained facial features that complement the temporal representations derived from video data. Dual-modal pre-training, which combines image (AffectNet) and video (VoxCeleb2) datasets (line 5), achieves the best performance, with a UAR/WAR of 43.41%/53.65% on FERV39K and 70.58%/79.37% UAR/WAR on DFEW. This confirms that dual-modal pre-training effectively captures both temporal information and fine-grained facial details, leading to robust representation capabilities. Last experiment in line 5, which utilized homogeneous image and video data from VoxCeleb2 for dual-modal pre-training, exhibited slightly lower performance than experiment in line 5. This suggests that leveraging diverse datasets and modalities for pre-training can yield richer and more beneficial feature representations. Overall, dual-modal pre-training significantly enhances DFER performance, surpassing both single-modal and no pre-training approaches.
TABLE IV: Comparison of the proposed dual-modal pre-training with other pre-training settings. Dual-modal pre-training on AffectNet and VoxCeleb2 significantly improves DFER performance and outperforms single-modal pre-training settings with less pre-training cost. AFC: AffectNet; VC2: VoxCeleb2; : Equal size with AffectNet.
Pre-Training Setting
(images + videos)
Pre-Training Cost
FERV39K DFEW
@@
Without pre-training 0 27.67 39.56 40.96 52.65
Image only (AFC) $8.8\times 10^{4}$ 36.10 47.59 53.80 66.82
Video only (VC2) $4.6\times 10^{5}$ 41.11 51.78 65.92 76.41
Video only (VC2) $1.8\times 10^{6}$ 42.92 52.58 67.71 76.33
Dual-modal (AFC + VC2) $\bm{1.4\times 10^{6}}$ 43.41 53.65 70.58 79.37
Dual-modal (VC2 + VC2) $1.4\times 10^{6}$ 42.36 52.94 68.35 78.72
@@ IV-D2 Pre-Training Cost Analysis
As shown in Table IV, the proposed dual-modal pre-training (line 5) demonstrates superior efficiency and effectiveness, significantly reducing computational costs compared to video-only pre-training (line 3). Specifically, the integration of image data reduces the pre-training cost by approximately 22% (from 1.8 to 1.4 $\times$ 106 TFLOPs), achieved through aggressive masking ratios of 90% for images and 95% for videos, minimizing redundant computations. This approach not only optimizes efficiency but also enhances performance, yielding notable improvements on the FERV39K (+0.49% UAR, +1.07% WAR) and DFEW (+2.87% UAR, +3.04% WAR) benchmarks. In contrast, image-only pre-training (line 2) incurs a minimal cost (8.8 $\times$ 104 TFLOPs) but shows limited effectiveness for DFER tasks. Similarly, the no pre-training setting (line 1) incurs no pre-training cost but severely degrades performance. These findings underscore the critical importance of capturing temporal dynamics for DFER, which are absent in image-only pre-training or no pre-training strategies.
@@
Model Setting Params (M) FERV39K DFEW
@@ IV-D3 Evaluation of the MoAE Module
To evaluate MoAE’s effectiveness in mitigating negative transfer between SFER and DFER tasks, we conducted comprehensive comparisons with other multi-task learning (MTL) methods. As shown in Table V, the vanilla ViT model achieves a performance of 42.68%/53.08% UAR/WAR on FERV39K and 68.69%/77.01% UAR/WAR on DFEW dataset. When SFER data is incorporated through simple multi-task learning (simply adding multiple task-specific heads to ViT, line 3), we observe minimal improvement or even slight performance degradation compared to the vanilla ViT. This phenomenon suggests that straightforward multi-task adaptation may fail to capture task relationships and lead to negative transfer. In contrast, our proposed S4D (with MoAE) demonstrates significant improvements across all metrics while maintaining reasonable parameter efficiency. Specifically, S4D achieves the best performance with 43.41%/53.65% UAR/WAR on FERV39K and 70.58%/79.37% UAR/WAR on DFEW, validating its effectiveness in learning both task-specific knowledge and generic features. To further investigate the architectural benefits, we compared S4D with ViT-MoE [46], a classic MoE architecture that replaces the standard FFN with FFN experts. Despite utilizing substantially more parameters (115M/285M vs. 90M/101M for testing/training), ViT-MoE exhibits poorer performance. These results emphasize that S4D’s superior performance stems from its architectural innovations rather than mere parameter scaling, highlighting its parameter efficiency and effectiveness in learning complementary knowledge between SFER and DFER tasks while mitigating negative transfer.
TABLE VI: Ablation studies on the joint fine-tuning strategy. Joint fine-tuning improves DFER significantly without notably compromising SFER. *: Results obtained at epoch 43, which yields optimal SFER performance. Other configurations are selected based on best DFER metrics.
Fine-Tuning Setting
(data mixture)
FERV39K AffectNet-7
@@ IV-D5 Contribution of Proposed Components
We conducted ablation studies to evaluate the impact of each component in our S4D framework: dual-modal pre-training (D.P.), MoAE, and joint fine-tuning (J.F.). Results from Table VII indicate that D.P. improved UAR/WAR by 0.10%/0.74% on FERV39K and 4.03%/0.85% on DFEW over the baseline (our fine-tuned version of VideoMAE on the pre-trained weights from [6]), underscoring its role in enhancing spatiotemporal representations. However, MoAE alone led to a slight performance decrease on all metrics, suggesting it may require a joint fine-tuning strategy to unlock its task-specific capabilities fully. The optimal configuration, integrating all three components (dual-modal pre-training, joint fine-tuning, and the MoAE module), achieved UAR/WAR of 43.41%/53.65% on FERV39K and 70.58%/79.37% on DFEW. These results highlight the synergistic effect of the full S4D framework, where the integration of data diversity, task coupling, and architectural enhancement collectively drives substantial improvements in DFER performance.